How do I solve: √96/(x√23) = 1/2
Answer by stanbon(75887)

√96/(x√23) = 1/2
Square both sides to get:
96/23x^2 = 1/4
23x^2 = 4*96
x^2 = 16.6957
x = +/-4.08604....
Cheers,
Stan H.